This triangle and square have the same perimeter.

Show that the square has an area 50% greater than the triangle.


Show all your working out.


1
Expert's answer
2021-02-24T06:32:39-0500

Let "a=" the side of the rectangle, "b=" the side of the equilateral triangle.

Given that triangle and square have the same perimeter


"3b=4a"

The area of the equilateral triangle


"S_1=\dfrac{\sqrt{3}}{4}b^2"

The area of the square


"S_2=a^2"

"\dfrac{S_2}{S_1}=\dfrac{a^2}{\dfrac{\sqrt{3}}{4}b^2}=\dfrac{4\sqrt{3}}{3}(\dfrac{3}{4})^2=\dfrac{3\sqrt{3}}{4}"


The square has an area approximately 30% greater than the equilateral triangle.